OpenStudy (anonymous):

Solve the following inequalities? Graph the inequality answers on a number line graph. 5p -18 = -43?? 7

OpenStudy (ivettef365):

first add 18 on both sides and you have 5p -18 +18 = -43 +18 5p = -25

OpenStudy (ivettef365):

now divide by 5, what do you get

OpenStudy (ivettef365):

and this is not an inequality ???

OpenStudy (anonymous):

it was 5p over 7 @ivettef365

OpenStudy (ivettef365):

oh, ok so the problem is 5/7p -18 = -43

OpenStudy (anonymous):

yes @ivettef365

OpenStudy (ivettef365):

ok so you start the same way 5/7p -18 +18 = -43 +18 5/7p = -25

OpenStudy (ivettef365):

now multiply by 7/5 so you can isolate p (7/5)(5/7) = 1 so p = -25(7/5) p = -5 (7) p = -35